OpenStudy (anonymous):

Find the standard form or the equation of the ellipse w/ vertices: (+-7,0) and eccentricity e=2/7

OpenStudy (amistre64):

id have to look up what eccentricity is lol

OpenStudy (amistre64):

x^2 y^2 --- + --- 49 b^2 is all i got without that bit of info

OpenStudy (amistre64):

eccentricity = foci point/ vertex

OpenStudy (anonymous):

ok that's helpful to know lol

OpenStudy (amistre64):

so the foci are at 2; and the vertex = 7 soo b^2 = 49 + 2^2

OpenStudy (anonymous):

wait wouldn't it be -

OpenStudy (amistre64):

sqrt(53) = b^2 so 53 should be tucked under the y right? which seems odd to me

OpenStudy (amistre64):

-2^2 yeaah, your right lol

OpenStudy (amistre64):

that dont seem so wierd to me now lol

OpenStudy (amistre64):

45 under the y is better lol

OpenStudy (anonymous):

wew ok I was thinking the same thing thanks so much I'm starting to get the hang of this silly math nonsense lol